OpenStudy (kainui):

What other kinds of identities exist like this for integers? cos(n pi)=(-1)^n

OpenStudy (kainui):

\[\cos(n \pi)=(-1)^n \]

OpenStudy (aravindg):

\[\sin n \pi=0\]

geerky42 (geerky42):

and \(\sin\left(\dfrac{n\pi}{2}\right) = (-1)^{n+1}\)

geerky42 (geerky42):

exactly what you are looking for?

OpenStudy (kainui):

Just anything fun or interesting @geerky42. Maybe some factorial stuff too. =P

OpenStudy (anonymous):

these identities are so beautiful

OpenStudy (kirbykirby):

\(\large e^{\pi i (2k+1)}=-1, k \in \mathbb{Z}\) ?

OpenStudy (kirbykirby):

not sure if that's what you're looking for

OpenStudy (kainui):

Yeah, for the most part. I don't know if that is really that much different than what's already been said though @kirbykirby